Average is 72, scored 53. What is the lowest scores can make to pass my exam with an average of 65?

Answer by JulietG(1812)

It depends upon the weight of the scores. If your 72 average is a third of your grade, the 53 another third, and the exam another third, it works this way:
Total points available = 300
Score desired = 65*3, or 195
Points already earned = 72+53, or 128
Points remaining to be scored for a 65 average = 195-128, or 67